(1 point) Given that ū = and are eigenvectors of the matrix -12 24 determine the corresponding eigenvalues. 21 = -1 12 = 1(1 point) Solve the system -6 1 dx dt х -6 -1 with the initial value 0 x(0) = -2 x(t) =(1 point) Calculate the eigenvalues of this matrix: [Note-- youll probably want to use a calculator or computer to estimate(1 point) For the linear system = [- :): Find the eigenvalues and eigenvectors for the coefficient matrix. 21 and 12 = V1 = V
Answer #1

* Given that ü [-] and - 2 are eigen vectors of the matrix - 12 A = 24 2 1-4 16 6 Now, The characteristic equation is a det (The given system of differential equation is 1 du da dt 6 -1 1. xa → dxn dt d22 ] let. x= 6 9 > dra dt - 6x1 + x2 & drz dt -Now, from X2 (D+ 6) * we get ONT x2 (8+6) (Cs e 3t + Ge4t) (-36 e 3t_ 4 Ce4t Gett) (3cje 3+ + 2 cecht). -3t +6cae the X2 x2 2matrix is The given I 16 4o A= -80 - 32 Then, the characteristic equation is det (A - I) 16- X 40 1- - 32 -80- ² t > (16-2) (Then, (A - 2 1. Iz Iz) y = 0 + 80 > 40 ya - 0 (zero matrix) -32 - 16 Y2 mon a 804, +40 y = 0 ) - 3241 164, = 0 Therefore, 2 yThe * given linear system is ģ = [34] y 1 ody, at dyz dt 6 The 3 -4 on efficient matrix is A= AI) = 0 Then, A 5- 6 -3 - 4-0be the eigenvector to 2 2 Again, let 3 [5. corresponding (A - AIZ) = 0 66] [] Then 3 > 3 त्र 34 + 6Y2 = 0 & -34, -64₂=0 7 + y